Arlington Clothing, Inc., shows the following information for its two divisions for year 1.

Lake Region Coastal Region
Sales revenue $ 4,200,000 $ 13,110,000
Cost of sales 2,711,300 6,555,000
Allocated corporate overhead 252,000 786,600
Other general and administration 557,900 3,759,000


Required:

a. Compute divisional operating income for the two divisions. Ignore taxes.

b-1. What are the gross margin and operating margin percentages for both divisions?

b-2. How well have these divisions performed?
Answer #1
Lake Region Coastal Region
i Sales revenue 4,200,000 13,110,000
ii Cost of sales 2,711,300 6,555,000
iii Allocated corporate overhead 252,000 786,600
iv Other general and administration 557,900 3,759,000
ans a
v=i-ii-iii-iv Operating income = 678,800 2,009,400
ans b-1
vii=i-ii Gross margin 1,488,700 6,555,000
viii=vii/i Gross margin % 35.45% 50.00%
vi=v/i Operating income %= 16.16% 15.33%
ans b-2 From above we can see that gross margin of Coastal region is higher
but operating income % of Lake region higher.
this is because general and administration overhead for Coaster region is very high.